OpenStudy (madeleine2):

p(x) is a polynomial with integer coefficients and p(-3) = 0. Which statements must be true? Choose all that apply. a. x + 3 is a factor of the polynomial. b. x - 3 is a factor of the polynomial. c. -3 is the constant term of the polynomial. d. p(x) can have at most 3 linear factors.

satellite73 (satellite73):

for example if \[(x+5)(x+3)=0\]the one solution is \(x=-3\)

OpenStudy (josh0404):

the answer is A. because when you equate p(x) is (-3) x+3=0 (-3)+3=0 0=0